# Mellin‑Space Prony Representability of Linear Viscoelastic Models

### v1

@@ 5.2 Coalescent Poles Cases
    If any spacing $1/c_{j}$ or $1/\alpha$ exceeds 1, by Lemma 1the corresponding progression cannot be a subset of the integers, and in fact has gaps of size at least 1; consequently, the union of such progressions leaves infinitely many integers uncovered.
@@
Since the necessary cancellations and residue compatibilities cannot be satisfied simultaneously for generic parameters, no $\widetilde{H}(s)\in\mathcal{P}$ can satisfy (15). Hence the constitutive equation admits no solution within the finite Prony class $\mathcal{P}$. ∎
@@ 6 Classification of Common Viscoelastic Models
$\Delta_{G}$: spacing of the dominant Gamma‑factor pole lattice. Transcendental models require infinite Prony ladders (Theorem 2). Notation: $0<\alpha,\beta<1$, $\Delta G=G_{0}-G_{\infty}$, $G_{e}$ = equilibrium modulus, $\delta=G_{e}/G_{\infty}<1$. The Cole–Cole entry uses the viscoelastic sign convention.
@@
$\tilde{G}(s)=G_{0}\tau_{0}^{\beta}\delta(s-\beta)$ is a Dirac delta, which cannot be expressed as $K(s)\tilde{H}(-s)$ with $\tilde{H}\in\mathcal{P}$ because the right-hand side has poles on integer lattices (from $K(s)$) and possibly other arithmetic progressions, not a single isolated pole.
@@ 8 Conclusions
The Mellin‑space framework unifies the treatment of power‑law, Gaussian, Cole–Cole, Havriliak–Negami, Cole–Davidson, and Mittag‑Leffler models, revealing a direct correspondence between fractional‑order operators and the arithmetic geometry of their kernel singularities.
@@ A.1 Derivation of the Mellin-space constitutive equation
For $\tau>0$ and $0<\Re s<1$, the Mellin transform (in $\omega$) of $\frac{i\omega\tau}{1+i\omega\tau}$ is given by the meromorphic continuation
@@ Appendix B Auxiliary Lemmas for the Proof of Theorem 1
Thus the only solution to the coupled recurrence is the trivial one. For a non‑trivial $\tilde{H}(s)$ to exist, the recurrence must decouple: each aligned sublattice must involve residues from only one pole family. ∎

### v2

@@ 5.2 Coalescent Poles Cases
    If any spacing $1/c_{j}$ or $1/\alpha$ is not an integer, the corresponding progression cannot be a subset of the set of integers. Even if the spacing equals an integer $q>1$, the progression contains only every $q$-th integer, leaving infinitely many integers uncovered.
@@
While general exact covers of the integers can exist with distinct moduli greater than 1, the specific progressions $\Lambda_{j}$ and $\Lambda_{\alpha}$ in our system originate from the distinct roots of the trial state and candidate Gamma factors. For generic, physical material parameters, these progressions do not possess the highly fine-tuned, commensurable rational dependencies required to form an exact congruence system that tiles $\mathbb{Z}$. Therefore, the required covering cannot be achieved.
@@
Since the necessary cancellations and residue compatibilities cannot be satisfied simultaneously for generic parameters, no $\widetilde{H}(s)\in\mathcal{P}$ can satisfy (15). Hence the constitutive equation admits no solution within the finite Prony class $\mathcal{P}$.
@@ 6 Classification of Common Viscoelastic Models
$\Delta_{G}$: spacing of the dominant Gamma‑factor pole lattice. Transcendental models require infinite Prony ladders (Theorem 2). Notation: $0<\alpha,\beta<1$, $\Delta G=G_{0}-G_{\infty}$, $G_{e}$ = equilibrium modulus, $\delta=G_{e}/G_{\infty}<1$. The Cole–Cole entry uses the viscoelastic sign convention. The Gaussian spectrum has no closed-form $G^{*}(\omega)$ but is included as a canonical transcendental example.
@@
$\tilde{G}(s)=G_{0}\tau_{0}^{\beta}\delta(s-\beta)$ is a Dirac delta, which cannot be expressed as $K(s)\tilde{H}(-s)$ with $\tilde{H}\in\mathcal{P}$ because the right-hand side has poles on integer lattices (from $K(s)$) and possibly other arithmetic progressions, not a single isolated pole. ∎
@@ 7 Test for Finite Prony Representability
Proposition 6 (Log-normal spectrum is transcendental).
The log-normal relaxation spectrum ( i.e. Gaussian in $\ln\tau$-variable)
$H_{\mathrm{LN}}(\tau)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(\ln\tau-\mu)^{2}}{2\sigma^{2}}\right),\qquad\tau>0,$
has Mellin transform
$\tilde{H}_{\mathrm{LN}}(s)=\exp\left(\mu s+\frac{\sigma^{2}}{2}s^{2}\right),$
which is entire of order 2. Consequently, $H_{\mathrm{LN}}(\tau)$ is not in the finite Prony class $\mathcal{P}$ and therefore belongs to the transcendental class.
Proof.
The Mellin transform is computed via the substitution $u=\ln\tau$:
$\tilde{H}_{\mathrm{LN}}(s)=\int_{0}^{\infty}H_{\mathrm{LN}}(\tau)\tau^{s-1}d\tau=\int_{-\infty}^{\infty}\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-(u-\mu)^{2}/(2\sigma^{2})}e^{(s-1)u}e^{u}du.$
Completing the square in the exponent,
$(u-\mu)^{2}/(2\sigma^{2})+su=-\frac{[u-(\mu+\sigma^{2}s)]^{2}}{2\sigma^{2}}+\mu s+\frac{\sigma^{2}}{2}s^{2}.$
The Gaussian integral evaluates to $\sqrt{2\pi\sigma^{2}}$, yielding the stated entire function.
By Corollary 1, any Mellin transform that is entire and not of the affine form $e^{\alpha s+\beta}$ cannot be a finite exponential sum. Here $\tilde{H}_{\mathrm{LN}}(s)$ is entire of order 2 ($\sigma^{2}>0$), hence not affine. Proposition 1 then implies $H_{\mathrm{LN}}(\tau)$ is not a finite Prony series.
The infinite ladder construction follows from Theorem 2, with distributional convergence guaranteed by the Mellin-Poisson summation formula [13]. ∎
Remark 5.
Recent work by Uneyama [14] used information-theoretic arguments to show that the log-normal distribution is the most probable relaxation spectrum under minimal assumptions. Proposition 6 reveals that this ”most probable” spectrum is fundamentally transcendental within our classification—it requires an infinite Prony ladder for exact representation. This synthesis suggests a powerful duality: information theory identifies the generic spectral shape, while pole-lattice analysis explains why it cannot be realized by any finite mechanical network.
@@ 8 Conclusions
The Mellin‑space framework unifies the treatment of power‑law, Gaussian, log-normal Cole–Cole, Havriliak–Negami, Cole–Davidson, and Mittag‑Leffler models, revealing a direct correspondence between fractional‑order operators and the arithmetic geometry of their kernel singularities.
@@ A.1 Derivation of the Mellin-space constitutive equation
For $\tau>0$ and $1<\Re s<0$, the Mellin transform (in $\omega$) of $\frac{i\omega\tau}{1+i\omega\tau}$ is given by the meromorphic continuation
@@ Appendix B Auxiliary Lemmas for the Proof of Theorem 1
Thus the only solution to the coupled recurrence is the trivial one. For a non‑trivial $\tilde{H}(s)$ to exist, the recurrence must decouple: each aligned sublattice must involve residues from only one pole family.
Asymptotic analysis.
Consider the ratio $r_{n}=R_{j,k_{n}}/R_{k,\ell_{n}}$. Using Gamma-residue formulas and Stirling’s approximation:
$r_{n}\sim\frac{C_{j}}{C_{k}}\frac{\Gamma(\ell_{n}+1)}{\Gamma(k_{n}+1)}.$
Then two cases arise:
Case 1: $a\neq c$. Then $\ell_{n}/k_{n}\to c/a\neq 1$, and
$r_{n}\sim Cn^{(c-a)n}e^{(a-c)n},$
which exhibits super-exponential growth or decay. Equation (B.3) divided by the larger residue then forces the dominant term to match $\rho_{n}$, impossible for generic parameters.
Case 2: $a=c$. Here $\ell_{n}=an+d$, $k_{n}=an+b$, so $\ell_{n}-k_{n}=d-b$ is constant. Stirling gives polynomial scaling:
$r_{n}\sim\frac{C_{j}}{C_{k}}n^{d-b}.$
All three sequences $R_{j,k_{n}}$, $R_{k,\ell_{n}}$, and $\rho_{n}$ now share the same factorial decay $1/(an)!$, but with polynomial prefactors $n^{b}$, $n^{d}$, and $A(s_{n})$ respectively. For generic parameters, these prefactors are incommensurate, and the system (B.3) remains infinitely overdetermined. The only solution is the trivial one.
Thus in both cases, the only solution to the coupled recurrence is $R_{j}=R_{k}=0$, implying $\tilde{H}\equiv 0$.